OpenStudy (anonymous):

Can anyone help me!?!?!?!?! Here is my question: Compare the functions shown below: f(x) = 4 sin (2x − π) - 1 g(x) x y -1 6 0 1 1 -2 2 -3 3 -2 4 1 5 6 h(x) = (x − 2)2 + 4 Which function has the smallest minimum y-value?

OpenStudy (anonymous):

can you help me @nelsonjedi

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the minimum value of \(f(x) = 4 \sin (2x − π) - 1 \) is if sine is equal to \(-1\) in which case you would get \[4\times (-1)-1=-5\]

OpenStudy (anonymous):

Smallest minimum value What does that mean? I could read it either way?

OpenStudy (anonymous):

the minimum value of \(g\) is apparently \(-3\)

OpenStudy (anonymous):

and the minimum value of \(h(x) = (x − 2)^2 + 4 \) is \(4\) since \((x-2)^2\geq 0\) always

OpenStudy (anonymous):

so whats the the lowest for f(x)

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

my best answer is above

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

can you help me @kabrona

OpenStudy (anonymous):

i'll give you a medal and be your fan

OpenStudy (anonymous):

or nah

OpenStudy (anonymous):

anyways look up the song nae nae

OpenStudy (anonymous):

LoL @maemaehoowahh

OpenStudy (anonymous):

The correct answer was f(x), just took the test :)

OpenStudy (anonymous):

@malice you're a saint. thank you :D